OpenStudy (megannicole51):

Use comparison test to determine convergence The series of ((2^(n)+1)/(n2^(n)-1)) n=1 b=infinity

OpenStudy (anonymous):

\[\sum \frac{2^n+1}{n(2^n-1)}\]?

OpenStudy (megannicole51):

the den.....n isnt outside the ( )

OpenStudy (anonymous):

\[\sum \frac{2^n+1}{n2^n-1}\]

OpenStudy (anonymous):

don't think it really makes any difference

OpenStudy (megannicole51):

well if u factor the n it would change to n2^n-n

OpenStudy (anonymous):

ok remember what series i said what the archetypical example of a divergent series even though the terms go to zero?

OpenStudy (megannicole51):

(1/n)?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so before we do any proof or anything, lets look at what we got

OpenStudy (anonymous):

\[\sum \frac{2^n+1}{2^n-1}\] certainly diverges, because the limit of the terms is not even 0, it is 1

OpenStudy (megannicole51):

Bn=(1/n) right? and i know what An is

OpenStudy (anonymous):

then if we put an \(n\) in the denominator and get \[\sum \frac{2^n+1}{n2^n-1}\] the terms do go to zero, but not fast enough because you only have the \(n\) down there

OpenStudy (anonymous):

this is my explanation not of how to show that it diverges, but how to know before you start that it diverges, so you can proceed now you have the job of proving it, and you are right, you can compare it to \(\frac{1}{n}\)

OpenStudy (anonymous):

and now i am off because i am on east coast time and it is late i am sure you can finish from there you will get used to these pretty fast i am sure

OpenStudy (megannicole51):

thank you so much for all of your help! seriously i want a good grade on this exam and so far i feel very confident...working through these practice problems with someone is such a great help!

OpenStudy (anonymous):

you are quite welcome good luck with your exam, i am sure you will do well

OpenStudy (megannicole51):

i have until wednesday to study so ill be on here a lot going over problems:) thank you